OpenStudy (anonymous):

9/8=k+6/6 answer with work?

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

Solving for K?

OpenStudy (anonymous):

Yeah i think, it says solve each proportion. Leave your answer as a fraction in simplest form

OpenStudy (anonymous):

Not sure how you need to write it out but the answer is 1/8

OpenStudy (anonymous):

If you isolate K by subtracting 6/6 from both sides you solve for K=1/8

OpenStudy (anonymous):

Where does the 9 go

OpenStudy (anonymous):

9/8=K + 6/6 -6/6 - 6/6

OpenStudy (anonymous):

Well in order to subtract 9/8 - 6/6 you need to change to common denominator of 24. So for subtraction you would turn the 9/8 into 27/24 and the 6/6 into 24/24. The common denominator for 8 and 6 is 24.

OpenStudy (anonymous):

Then 27/24-24/24=3/24 simplified into 1/8

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Sure thing